# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ 1.2. Quantum minors
(1.10) $u^{-1}\ldots(u-m\hbar+\hbar)^{-1}T_{b_{1}\ldots b_{m}}^{a_{1}\ldots a_{m}}(u)\in\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})[[u^{-1}]].$
@@
    $T_{b_{1}\ldots b_{m}}^{a_{\sigma(1)}\ldots a_{\sigma(m)}}(u)=\mathrm{sgn}(\sigma)T_{b_{1}\ldots b_{m}}^{a_{1}\ldots a_{m}}=T_{b_{\sigma(1)}\ldots b_{\sigma(m)}}^{a_{1}\ldots a_{m}}(u).$
    In particular, if $a_{i}=a_{j}$ or $b_{i}=b_{j}$ for some $i\neq j$, then $T_{b_{1}\ldots b_{m}}^{a_{1}\ldots a_{m}}(u)=0$.
@@
    [Mol07, Proposition 1.6.8] A quantum minor can be decomposed as
    $T_{b_{1}\ldots b_{m}}^{a_{1}\ldots a_{m}}(u)$ $=\sum\limits_{l=1}^{m}(-1)^{m-l}T_{b_{1}\ldots b_{m-1}}^{a_{1}\ldots\hat{a}_{l}\ldots a_{m}}(u)T_{a_{l}b_{m}}(u-\hbar m+\hbar)=$
    $=\sum\limits_{l=1}^{m}(-1)^{m-l}T_{b_{1}\ldots\hat{b}_{l}\ldots b_{m}}^{a_{1}\ldots a_{m-1}}(u-\hbar)T_{a_{m}b_{l}}(u)=$
    $=\sum\limits_{l=1}^{m}(-1)^{l-1}T_{a_{l}b_{1}}(u)T_{b_{2}\ldots b_{m}}^{a_{1}\ldots\hat{a}_{l}\ldots a_{m}}(u-\hbar)=$
    $=\sum\limits_{l=1}^{m}(-1)^{l-1}T_{a_{1}b_{l}}(u-\hbar m+\hbar)T_{b_{1}\ldots\hat{b}_{l}\ldots b_{m}}^{a_{2}\ldots a_{m}}(u).$
@@
    [Mol07, Proposition 1.7.1] The following relation holds:
    $E_{kl},T^{a_{1}\ldots a_{m}}_{b_{1}\ldots b_{m}}(u)]=\hbar\sum_{i=1}^{m}\delta_{a_{i}l}T^{a_{1}\ldots k\ldots a_{m}}_{b_{1}\ldots b_{m}}(u)-\hbar\sum_{i=1}^{m}\delta_{b_{i}k}T^{a_{1}\ldots a_{m}}_{b_{1}\ldots l\ldots b_{m}}(u),$
@@
For the last part, it is enough to consider the coefficient of $u^{0}$ on both sides of [Mol07, Proposition 1.7.1]. ∎
@@
(1.11) $A(u):=(-1)^{N}T^{1\ldots N}_{1\ldots N}(-u+N\hbar-\hbar)=:\sum\limits_{i=0}^{N}A_{i}u^{N-i}.$
By [MNO96], the coefficients $\{A_{i}\}$ generate the center of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$.
We will need an explicit form of the inverse $T(u)^{-1}$ in terms of quantum comatrix, see [Mol07, Proposition 1.9.2]:
(1.12) $\hat{T}_{ij}(u):=(-1)^{i+j}T_{1\ldots\hat{i}\ldots N}^{1\ldots\hat{j}\ldots N}(u)$
where the hats indicate the omitted indices. The following equation is an immediate corollary of [Mol07, Definition 1.9.1] and [Mol07, Proposition 1.9.2]:
(1.13) $T(u)^{-1}=(-1)^{N}\hat{T}(u+N\hbar-\hbar)A(-u)^{-1}.$
Recall the antiautomorphism (1.7). The next lemma follows immediately from the definition (1.9) and $\tilde{\varpi}_{N}[T(u)]=-T(-u)^{T}$, where the superscript means matrix transposition.
@@
1.3. Differential operators
Let $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ be the Rees algebra of the differential operators on $\mathrm{GL}_{N}$ with respect to the degree filtration. For $\xi\in\mathfrak{gl}_{N}$, consider the corresponding vectors fields generating the right (resp. the left) translations along $\xi$:
$\xi^{R}(f)(g):=\frac{\mathrm{d}}{\mathrm{d}t}f(ge^{t\xi})\Big{|}_{t=0},\qquad\xi^{L}(f)(g):=\frac{\mathrm{d}}{\mathrm{d}t}f(e^{-t\xi}g)\Big{|}_{t=0}.$
@@ 1.3. Differential operators
Denote by $\partial_{ij}:=\partial/\partial x_{ij}$ and by $D=(\partial_{ij})$ the corresponding matrix. We have
$E_{ij}^{R}=\sum_{\alpha=1}^{N}x_{\alpha i}\partial_{\alpha j},\qquad E_{ij}^{L}=-\sum_{\alpha=1}^{N}x_{j\alpha}\partial_{i\alpha}.$
In particular, in terms of matrices from Section˜1.1, we can write
$E^{R}=(E_{ij}^{R})=XD,\qquad E^{L}=(E_{ij}^{L})=-DX+\hbar N\cdot\mathrm{Id}_{N},$
@@
(1.14) $E^{L}=X^{-1}(-E^{R}+\hbar N\cdot\mathrm{Id}_{N})X.$
@@
It is enough to show that their actions on $\mathcal{O}(\mathrm{Mat}_{N\times N})$ coincide. By the Howe duality [CW12, Theorem 5.16], under the $(\mathfrak{gl}_{N},\mathfrak{gl}_{N})$-action by
$E^{(1)}_{ij}\mapsto\sum_{\alpha=1}^{N}x_{\alpha i}\partial_{\alpha j},\qquad E^{(2)}_{ij}\mapsto\sum_{\alpha=1}^{N}x_{i\alpha}\partial_{j\alpha},$
@@
an algebra automorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$, a variant of (1.6). Then $E^{(1)}_{ij}=E^{R}_{ij}$ and $E^{(2)}_{ij}=\tilde{\varpi}_{N}(E^{L}_{ij})$, therefore,
$A^{L}(u)=\tilde{\varpi}_{N}(A^{R}(u)).$
The result follows from ˜1.2. ∎
@@
It follows from Proposition˜7.2 and Proposition˜4.7 that
@@
where the equality is understood as in Remark˜4.8 and multiplication by $A^{R}(u)$ is entry-wise. Therefore, by (1.13) and $T^{R}(u)=u+E^{R}$, we have
$X^{-1}\cdot\hat{T}^{R}(-u+N\hbar-\hbar)\cdot X=(-1)^{N}X^{-1}\frac{1}{-u+E^{R}}A^{R}(u)X=(-1)^{N+1}X^{-1}\frac{1}{u-E^{R}+\hbar}XA^{R}(u).$
It follows from (1.14) and Proposition˜1.3 that
$(-1)^{N+1}X^{-1}\frac{1}{u-E^{R}+\hbar}XA^{R}(u)=-\frac{1}{u+E^{L}-\hbar N+\hbar}A^{L}(-u+N\hbar-\hbar)=$
@@
2. Degenerate affine Hecke algebras and Yangians
2.1. Shibukawa-Ueno $R$-operator
In this subsection, we recall the definition of Shibukawa-Ueno $R$-operators [SU92] and study the induced symmetrizers. By [FP94], its elliptic version is gauge-equivalent to Belavin’s $R$-matrix that is used to define the original elliptic shuffle algebras [OF89, OF97, FO98]. At the same time, we will show in the next subsection that its rational version is related to degenerate affine Hecke algebras.
Let $\theta(z)$ be a skew-symmetric (i.e. $\theta(-z)=-\theta(z)$) holomorphic function satisfying the three-term relation
(2.1) $\begin{split}\theta(x+&y)\theta(x-y)\theta(z+w)\theta(z-w)+\\ +&\theta(x+z)\theta(x-z)\theta(w+y)\theta(w-y)+\\ +&\theta(x+w)\theta(x-w)\theta(y+z)\theta(y-z)=0.\end{split}$
Its analytic solution is essentially one of the following three functions [WW96]:
(2.2) $\theta^{\mathrm{ell}}(z):=\vartheta_{1}(z,\tau),\qquad\theta^{\mathrm{trig}}(z):=\sin(\pi z/\tau_{1}),\qquad\theta^{\mathrm{rat}}(z):=z,$
@@ 2.1. Shibukawa-Ueno $R$-operator
$\vartheta_{1}(z,\tau)=-\sum_{m\in\mathbb{Z}}\exp\left[\pi i\tau\left(m+\frac{1}{2}\right)^{2}+2\pi i\left(m+\frac{1}{2}\right)\left(z+\frac{1}{2}\right)\right]$
@@
$G^{\theta}(z,w):=\frac{\theta(z+w)\theta^{\prime}(0)}{\theta(z)\theta(w)}$
Let $\mathcal{M}_{k}$ the space of meromorphic functions in variables $z_{1},\ldots,z_{k}$. Denote by
(2.3) $(P_{ij}f)(z_{1},\ldots,z_{i},\ldots,z_{j},\ldots,z_{k}):=f(z_{1},\ldots,z_{j},\ldots,z_{i},\ldots,z_{k})$
@@
Definition 2.1.
[SU92] The Shibukawa-Ueno $R$-operator associated to $\theta$ is the operator on $\mathcal{M}_{k}$ defined by
@@
The Shibukawa-Ueno braiding element is
@@
The elliptic (resp. trigonometric, rational) Shibukawa-Ueno $R$-operator is
@@
Proposition 2.2.
[SU92] The Shibukawa-Ueno $R$-operator satisfies the quantum Yang-Baxter equation
@@
Consider the $\kappa$-symmetrization operator [Che86]
$\check{H}_{k}^{\theta}(\kappa):=(\check{R}^{\theta}_{k-1}(-\kappa|\kappa)\ldots\check{R}^{\theta}_{1}(-k\kappa+\kappa|\kappa))\ldots(\check{R}^{\theta}_{k-1}(-\kappa|\kappa)\check{R}^{\theta}_{k-2}(-2\kappa|\kappa))(\check{R}^{\theta}_{k-1}(-\kappa|\kappa))$
@@
$(\mathrm{Sym}_{k}f)(z_{1},\ldots,z_{k}):=\sum_{\sigma\in S_{k}}f(z_{\sigma(1)},\ldots,z_{\sigma(k)})$
@@
$(\check{H}_{k}^{\theta}(\kappa)f)(z_{1},\ldots,z_{k})=\mathrm{Sym}_{k}\left(\prod_{i<j}G^{\theta}(z_{i}-z_{j},\kappa)f(z_{1},\ldots,z_{k})\right).$
@@
$H_{k}^{\theta}=(R_{12}^{\theta}(-\kappa)\ldots R_{1k}(-k\kappa+\kappa))\ldots(R^{\theta}_{k-2,k-1}(-\kappa)R^{\theta}_{k-2,k}(-2\kappa))(R^{\theta}_{k-1,k}(-\kappa)),$
@@
$(H_{k}^{\theta}f)(z_{1},\ldots,z_{k})=\mathrm{Sym}_{k}\left(\prod_{i<j}G^{\theta}(z_{i}-z_{j},\kappa)f(z_{1},\ldots,z_{k})\right).$
@@
$F(z_{0},z_{1},z_{2},\ldots,z_{k}):=(H_{k}^{\theta}f)(z_{0},z_{1},\ldots,z_{k})=\mathrm{Sym}_{k}\left(\prod_{1\leq i<j\leq k}G^{\theta}(z_{i}-z_{j})f(z_{0},z_{1},\ldots,z_{k})\right),$
where symmetrization is over the variables $z_{1},\ldots,z_{k}$; in particular, $F$ is symmetric with respect to them. Then it would be enough to prove that
$R_{01}(-\kappa)\ldots R_{0k}(-k\kappa)F=(1+P_{01}+\ldots+P_{0k})\prod_{i=1}^{k}G^{\theta}(z_{0}-z_{i},\kappa)F(z_{0},\ldots,z_{k}).$
Denote by $\mathcal{M}_{k+1}^{S_{l}}$ the space of meromorphic functions in $z_{0},z_{1},\ldots,z_{k}$ symmetric with respect to the variables $z_{1},\ldots,z_{l}$. For any $1\leq l\leq k$, define
$X_{l}\colon\mathcal{M}_{k+1}^{S_{l}}\rightarrow\mathcal{M}_{k+1}^{S_{l}},\qquad X_{l}=R^{\theta}_{01}(-\kappa)\ldots R_{0l}(-l\kappa).$
@@
$X_{l}=(1+P_{01}+\ldots+P_{0l})\prod_{i=1}^{l}G^{\theta}(z_{0}-z_{i}).$
For $l=1$ it follows from $\theta(-z)=-\theta(z)$ for (2.2). For the step, we have
$X_{l+1}$ $=X_{l}R_{0,l+1}(-l\kappa-\kappa)=$
$=(1+P_{01}+\ldots+P_{0l})\prod_{i=1}^{l}G^{\theta}(z_{0}-z_{i},\kappa)\cdot(G^{\theta}(z_{0}-z_{l+1},\kappa)+G^{\theta}(z_{l+1}-z_{0},l\kappa+\kappa)P_{0,l+1})$
@@
$(1+P_{01}+\ldots+P_{0l})\prod_{i=1}^{l}G^{\theta}(z_{0}-z_{i},\kappa)\cdot G^{\theta}(z_{l+1}-z_{0},l\kappa+\kappa)P_{0,l+1}=P_{0,l+1}\prod_{i=1}^{l+1}G^{\theta}(z_{0}-z_{i},\kappa).$
The right-hand side is $\prod_{i=0}^{l}G^{\theta}(z_{l+1}-z_{i},\kappa)P_{0,l+1}$. Observe that we have $P_{0i}P_{0,l+1}=P_{0,l+1}$ for any $1\leq i\leq l$ as operators on $\mathcal{M}_{k+1}^{S_{l+1}}$. Therefore, the left-hand side is
$\sum_{i=0}^{l}\prod_{\begin{subarray}{c}j=0\ldots l\\ j\neq i\end{subarray}}G^{\theta}(z_{i}-z_{j},\kappa)\cdot G^{\theta}(z_{l+1}-z_{i},l\kappa+\kappa)P_{0,l+1}.$
@@
$\prod_{i=0}^{l}G^{\theta}(z_{l+1}-z_{i},\kappa)=\sum_{i=0}^{l}\prod_{\begin{subarray}{c}j=0\ldots l\\ j\neq i\end{subarray}}G^{\theta}(z_{i}-z_{j},\kappa)\cdot G^{\theta}(z_{l+1}-z_{i},l\kappa+\kappa)$
This is the content of ˜A.1 for $u=z_{l+1}$. ∎
@@
$\check{A}^{\theta}_{k}(\kappa):=(\check{R}^{\theta}_{k-1}(\kappa|\kappa))(\check{R}^{\theta}_{k-2}(2\kappa|\kappa)\check{R}^{\theta}_{k-1}(\kappa|\kappa))\ldots(\check{R}^{\theta}_{1}(k\kappa-\kappa|\kappa)\ldots\check{R}^{\theta}_{k-1}(\kappa|\kappa))$
@@
$(\mathrm{Sym}^{-}_{k}f)(z_{1},\ldots,z_{k}):=\sum_{\sigma\in S_{k}}(-1)^{\mathrm{sgn}(\sigma)}f(z_{\sigma(1)},\ldots,z_{\sigma(k)})$
@@
$(\check{A}^{\theta}_{k}f)(z_{1},\ldots,z_{k})=\prod_{i<j}G^{\theta}(z_{i}-z_{j},-\kappa)\cdot\mathrm{Sym}^{-}_{k}f(z_{1},\ldots,z_{k}).$

### v4

@@ 2.2 Quantum minors
$u^{-1}\cdots(u-m\hbar+\hbar)^{-1}T_{b_{1}\dots b_{m}}^{a_{1}\dots a_{m}}(u)\in\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\bigl[\hskip-0.56905pt\bigl[u^{-1}\bigr]\hskip-0.56905pt\bigr].$
@@ 1. $(1)$
    $T_{b_{1}\dots b_{m}}^{a_{\sigma(1)}\dots a_{\sigma(m)}}(u)=\mathrm{sgn}(\sigma)T_{b_{1}\dots b_{m}}^{a_{1}\dots a_{m}}=T_{b_{\sigma(1)}\dots b_{\sigma(m)}}^{a_{1}\dots a_{m}}(u).$
    In particular, if $a_{i}=a_{j}$ or $b_{i}=b_{j}$ for some $i\neq j$, then $T_{b_{1}\dots b_{m}}^{a_{1}\dots a_{m}}(u)=0$.
@@ 2. $(2)$
    A quantum minor can be decomposed as [38, Proposition 1.6.8]
    $T_{b_{1}\dots b_{m}}^{a_{1}\dots a_{m}}(u)$ $=\sum\limits_{l=1}^{m}(-1)^{m-l}T_{b_{1}\dots b_{m-1}}^{a_{1}\dots\hat{a}_{l}\dots a_{m}}(u)T_{a_{l}b_{m}}(u-\hbar m+\hbar)$
    $=\sum\limits_{l=1}^{m}(-1)^{m-l}T_{b_{1}\dots\hat{b}_{l}\dots b_{m}}^{a_{1}\dots a_{m-1}}(u-\hbar)T_{a_{m}b_{l}}(u)$
    $=\sum\limits_{l=1}^{m}(-1)^{l-1}T_{a_{l}b_{1}}(u)T_{b_{2}\dots b_{m}}^{a_{1}\dots\hat{a}_{l}\dots a_{m}}(u-\hbar)$
    $=\sum\limits_{l=1}^{m}(-1)^{l-1}T_{a_{1}b_{l}}(u-\hbar m+\hbar)T_{b_{1}\dots\hat{b}_{l}\dots b_{m}}^{a_{2}\dots a_{m}}(u).$
@@ 3. $(3)$
    The following relation holds [38, Proposition 1.7.1]:
    $\bigl[E_{kl},T^{a_{1}\dots a_{m}}_{b_{1}\dots b_{m}}(u)\bigr]=\hbar\sum_{i=1}^{m}\delta_{a_{i}l}T^{a_{1}\dots k\dots a_{m}}_{b_{1}\dots b_{m}}(u)-\hbar\sum_{i=1}^{m}\delta_{b_{i}k}T^{a_{1}\dots a_{m}}_{b_{1}\dots l\dots b_{m}}(u),$
@@
For the last part, it is enough to consider the coefficient of $u^{0}$ on both sides of [38, Proposition 1.7.1]. ∎
@@
$A(u):=(-1)^{N}T^{1\dots N}_{1\dots N}(-u+N\hbar-\hbar)=:\sum\limits_{i=0}^{N}A_{i}u^{N-i}.$ (2.8)
By [39], the coefficients $\{A_{i}\}$ generate the center of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$.
We will need an explicit form of the inverse $T(u)^{-1}$ in terms of quantum comatrix, see [38, Proposition 1.9.2]:
$\hat{T}_{ij}(u):=(-1)^{i+j}T_{1\dots\hat{i}\dots N}^{1\dots\hat{j}\dots N}(u),$ (2.9)
where the hats indicate the omitted indices. The following equation is an immediate corollary of [38, Definition 1.9.1 and Proposition 1.9.2]:
$T(u)^{-1}=(-1)^{N}\hat{T}(u+N\hbar-\hbar)A(-u)^{-1}.$ (2.10)
Recall the antiautomorphism (2.6). The next lemma follows immediately from the definition (2.7) and $\tilde{\varpi}_{N}[T(u)]=-T(-u)^{\mathsf{T}}$, where the superscript means matrix transposition.
@@
2.3 Differential operators
Let $\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ be the Rees algebra of the differential operators on $\mathrm{GL}_{N}$ with respect to the degree filtration. For $\xi\in\mathfrak{gl}_{N}$, consider the corresponding asymptotic vectors fields generating the right (resp. the left) translations along $\xi$:
$\xi^{R}(f)(g):=\hbar\frac{\mathrm{d}}{\mathrm{d}t}f\bigl(g{\rm e}^{t\xi}\bigr)\Big|_{t=0},\qquad\xi^{L}(f)(g):=\hbar\frac{\mathrm{d}}{\mathrm{d}t}f\bigl({\rm e}^{-t\xi}g\bigr)\Big|_{t=0}.$
@@ 2.3 Differential operators
Denote by $p_{ij}:=\hbar\partial/\partial x_{ij}$ and by $D=(p_{ij})$ the corresponding matrix. We have
$E_{ij}^{R}=\sum_{\alpha=1}^{N}x_{\alpha i}p_{\alpha j},\qquad E_{ij}^{L}=-\sum_{\alpha=1}^{N}x_{j\alpha}p_{i\alpha}.$
In particular, in terms of matrices from Section 2.1, we can write
$E^{R}=\bigl(E_{ij}^{R}\bigr)=XD,\qquad E^{L}=\bigl(E_{ij}^{L}\bigr)=-DX+\hbar N\cdot\mathrm{Id}_{N},$
@@
$E^{L}=X^{-1}\bigl(-E^{R}+\hbar N\cdot\mathrm{Id}_{N}\bigr)X.$ (2.11)
@@
It is enough to show that their actions on $\mathcal{O}(\mathrm{Mat}_{N\times N})$ coincide. By the Howe duality [7, Theorem 5.16], under the $(\mathfrak{gl}_{N},\mathfrak{gl}_{N})$-action by
$E^{(1)}_{ij}\mapsto\sum_{\alpha=1}^{N}x_{\alpha i}p_{\alpha j},\qquad E^{(2)}_{ij}\mapsto\sum_{\alpha=1}^{N}x_{i\alpha}p_{j\alpha},$
@@
an algebra automorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$, a variant of (2.5). Then $E^{(1)}_{ij}=E^{R}_{ij}$ and $E^{(2)}_{ij}=\tilde{\varpi}_{N}\bigl(E^{L}_{ij}\bigr)$, therefore,
$A^{L}(u)=\tilde{\varpi}_{N}\bigl(A^{R}(u)\bigr).$
The result follows from Lemma 2.2. ∎
@@
It follows from Propositions 8.3 and 5.7 that
@@
where the equality is understood as in Remark 5.8 and multiplication by $A^{R}(u)$ is entry-wise. Therefore, by (2.10) and $T^{R}(u)=u+E^{R}$, we have
$X^{-1}\cdot\hat{T}^{R}(-u+N\hbar-\hbar)\cdot X$ $=(-1)^{N}X^{-1}\frac{1}{-u+E^{R}}A^{R}(u)X$
$=(-1)^{N+1}X^{-1}\frac{1}{u-E^{R}+\hbar}XA^{R}(u).$
It follows from (2.11) and Proposition 2.3 that
$(-1)^{N+1}X^{-1}\frac{1}{u-E^{R}+\hbar}XA^{R}(u)$ $=-\frac{1}{u+E^{L}-\hbar N+\hbar}A^{L}(-u+N\hbar-\hbar)$
@@
3 Degenerate affine Hecke algebras and Yangians
The goal of this section is to establish a relation between rational Feigin–Odesskii shuffle algebras and degenerate affine Hecke algebras summarized in Theorem 3.11. Using, on the one hand, the action of latter on certain Harish-Chandra bimodules and their reductions as in Section 6, on the other hand, connection between the former and (truncated shifted) Yangians of $\mathfrak{sl}_{2}$ which we recall in Section 3.3, we apply the results of this section to prove the main result of the paper, Theorem 8.10. This relation is deduced through Shibukawa–Ueno $R$-operators whose definition and properties we recall in Section 3.1.
3.1 Shibukawa–Ueno $\boldsymbol{R}$-operator
In this subsection, we recall the definition of Shibukawa–Ueno $R$-operators [44] and study the induced symmetrizers. By [13], its elliptic version is gauge-equivalent to Belavin’s $R$-matrix that is used to define the original elliptic shuffle algebras [12, 40, 41]. At the same time, we will show in the next subsection that its rational version is related to degenerate affine Hecke algebras.
Let $\theta(z)$ be a skew-symmetric (i.e., $\theta(-z)=-\theta(z)$) holomorphic function satisfying the three-term relation
$\theta(x+y)\theta(x-y)\theta(z+w)\theta(z-w)+\theta(x+z)\theta(x-z)\theta(w+y)\theta(w-y)$
$\qquad{}+\theta(x+w)\theta(x-w)\theta(y+z)\theta(y-z)=0.$ (3.1)
Its analytic solution is essentially one of the following three functions [47]:
$\theta^{\mathrm{ell}}(z):=\vartheta_{1}(z,\tau),\qquad\theta^{\mathrm{trig}}(z):=\sin(\pi z/\tau_{1}),\qquad\theta^{\mathrm{rat}}(z):=z,$ (3.2)
@@ 3.1 Shibukawa–Ueno $\boldsymbol{R}$-operator
$\vartheta_{1}(z,\tau)=-\sum_{m\in\mathbb{Z}}\exp\Biggl[\pi{\rm i}\tau\biggl(m+\frac{1}{2}\biggr)^{2}+2\pi{\rm i}\biggl(m+\frac{1}{2}\biggr)\biggl(z+\frac{1}{2}\biggr)\Biggr]$
@@
$G^{\theta}(z,w):=\frac{\theta(z+w)\theta^{\prime}(0)}{\theta(z)\theta(w)}.$
Let $\mathcal{M}_{k}$ the space of meromorphic functions in variables $z_{1},\dots,z_{k}$. Denote by
$(P_{ij}f)(z_{1},\dots,z_{i},\dots,z_{j},\dots,z_{k}):=f(z_{1},\dots,z_{j},\dots,z_{i},\dots,z_{k})$ (3.3)
@@
Definition 3.1 ([44]).
The Shibukawa–Ueno $R$-operator associated to $\theta$ is the operator on $\mathcal{M}_{k}$ defined by
@@
The Shibukawa–Ueno braiding element is
@@
The elliptic (resp. trigonometric, rational) Shibukawa–Ueno $R$-operator is
@@
Proposition 3.2 ([44]).
The Shibukawa–Ueno $R$-operator satisfies the quantum Yang–Baxter equation
@@
Consider the $\kappa$-symmetrization operator [8]
$\check{H}_{k}^{\theta}(\kappa):=\bigl(\check{R}^{\theta}_{k-1}(-\kappa|\kappa)\cdots\check{R}^{\theta}_{1}(-k\kappa+\kappa|\kappa)\bigr)\cdots\bigl(\check{R}^{\theta}_{k-1}(-\kappa|\kappa)\check{R}^{\theta}_{k-2}(-2\kappa|\kappa)\bigr)\bigl(\check{R}^{\theta}_{k-1}(-\kappa|\kappa)\bigr).$
@@
$(\mathrm{Sym}_{k}f)(z_{1},\dots,z_{k}):=\sum_{\sigma\in S_{k}}f(z_{\sigma(1)},\dots,z_{\sigma(k)})$
@@
$\bigl(\check{H}_{k}^{\theta}(\kappa)f\bigr)(z_{1},\dots,z_{k})=\mathrm{Sym}_{k}\biggl(\prod_{i<j}G^{\theta}(z_{i}-z_{j},\kappa)f(z_{1},\dots,z_{k})\biggr).$
@@
$H_{k}^{\theta}=\bigl(R_{12}^{\theta}(-\kappa)\cdots R_{1k}(-k\kappa+\kappa)\bigr)\cdots\bigl(R^{\theta}_{k-2,k-1}(-\kappa)R^{\theta}_{k-2,k}(-2\kappa)\bigr)\bigl(R^{\theta}_{k-1,k}(-\kappa)\bigr).$
@@
$\bigl(H_{k}^{\theta}f\bigr)(z_{1},\dots,z_{k})=\mathrm{Sym}_{k}\biggl(\prod_{i<j}G^{\theta}(z_{i}-z_{j},\kappa)f(z_{1},\dots,z_{k})\biggr).$
@@
$F(z_{0},z_{1},z_{2},\dots,z_{k}):={}$ $\bigl(H_{k}^{\theta}f\bigr)(z_{0},z_{1},\dots,z_{k})$
$={}$ $\mathrm{Sym}_{k}\biggl(\prod_{1\leq i<j\leq k}G^{\theta}(z_{i}-z_{j})f(z_{0},z_{1},\dots,z_{k})\biggr),$
where symmetrization is over the variables $z_{1},\dots,z_{k}$; in particular, $F$ is symmetric with respect to them. Then it would be enough to prove that
$R_{01}(-\kappa)\cdots R_{0k}(-k\kappa)F=(1+P_{01}+\dots+P_{0k})\prod_{i=1}^{k}G^{\theta}(z_{0}-z_{i},\kappa)F(z_{0},\dots,z_{k}).$
Denote by $\mathcal{M}_{k+1}^{S_{l}}$ the space of meromorphic functions in $z_{0},z_{1},\dots,z_{k}$ symmetric with respect to the variables $z_{1},\dots,z_{l}$. For any $1\leq l\leq k$, define
$X_{l}\colon\ \mathcal{M}_{k+1}^{S_{l}}\rightarrow\mathcal{M}_{k+1}^{S_{l}},\qquad X_{l}=R^{\theta}_{01}(-\kappa)\cdots R_{0l}(-l\kappa).$
@@
$X_{l}=(1+P_{01}+\dots+P_{0l})\prod_{i=1}^{l}G^{\theta}(z_{0}-z_{i}).$
For $l=1$, it follows from $\theta(-z)=-\theta(z)$ for (3.2). For the step, we have
$X_{l+1}={}$ $X_{l}R_{0,l+1}(-l\kappa-\kappa)$
$={}$ $(1+P_{01}+\dots+P_{0l})$
${\times}\,\prod_{i=1}^{l}G^{\theta}(z_{0}-z_{i},\kappa)\cdot\bigl(G^{\theta}(z_{0}-z_{l+1},\kappa)+G^{\theta}(z_{l+1}-z_{0},l\kappa+\kappa)P_{0,l+1}\bigr).$
@@
$(1+P_{01}+\dots+P_{0l})\prod_{i=1}^{l}G^{\theta}(z_{0}-z_{i},\kappa)\cdot G^{\theta}(z_{l+1}-z_{0},l\kappa+\kappa)P_{0,l+1}$
$\qquad=P_{0,l+1}\prod_{i=1}^{l+1}G^{\theta}(z_{0}-z_{i},\kappa).$
The right-hand side is $\prod_{i=0}^{l}G^{\theta}(z_{l+1}-z_{i},\kappa)P_{0,l+1}$. Observe that we have $P_{0i}P_{0,l+1}=P_{0,l+1}$ for any $1\leq i\leq l$ as operators on $\mathcal{M}\raisebox{-0.5pt}{${}_{k+1}^{S_{l+1}}$}$. Therefore, the left-hand side is
$\sum_{i=0}^{l}\prod_{\begin{subarray}{c}j=0,\dots,l,\\ j\neq i\end{subarray}}G^{\theta}(z_{i}-z_{j},\kappa)\cdot G^{\theta}(z_{l+1}-z_{i},l\kappa+\kappa)P_{0,l+1}.$
@@
$\prod_{i=0}^{l}G^{\theta}(z_{l+1}-z_{i},\kappa)=\sum_{i=0}^{l}\prod_{\begin{subarray}{c}j=0,\dots,l,\\ j\neq i\end{subarray}}G^{\theta}(z_{i}-z_{j},\kappa)\cdot G^{\theta}(z_{l+1}-z_{i},l\kappa+\kappa).$
This is the content of Lemma A.1 for $u=z_{l+1}$. ∎
@@
$\check{A}^{\theta}_{k}(\kappa):=\bigl(\check{R}^{\theta}_{k-1}(\kappa|\kappa)\bigr)\bigl(\check{R}^{\theta}_{k-2}(2\kappa|\kappa)\check{R}^{\theta}_{k-1}(\kappa|\kappa)\bigr)\cdots\bigl(\check{R}^{\theta}_{1}(k\kappa-\kappa|\kappa)\cdots\check{R}^{\theta}_{k-1}(\kappa|\kappa)\bigr).$
@@
$\bigl(\mathrm{Sym}^{-}_{k}f\bigr)(z_{1},\dots,z_{k}):=\sum_{\sigma\in S_{k}}(-1)^{\mathrm{sgn}(\sigma)}f(z_{\sigma(1)},\dots,z_{\sigma(k)})$
@@
$\bigl(\check{A}^{\theta}_{k}f\bigr)(z_{1},\dots,z_{k})=\prod_{i<j}G^{\theta}(z_{i}-z_{j},-\kappa)\cdot\mathrm{Sym}^{-}_{k}f(z_{1},\dots,z_{k}).$